HUBBLE SPACE TELESCOPE DAILY REPORT #897 PERIOD COVERED: 0800 EST Daily Status Report as of 148/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4381 (HST Medium - Deep Survey Cycle 3) The Wide Field Camera was used to make one high latitude medium - deep sky survey image taken in parallel with the prime science observations. The observation was monitored in the OSS, and was reported to look fine. 1.2 Completed Two Sets of FOC/96 3801 (Search to QSOs Suitable for Subsequent Observations of He II) The Faint Object Camera was used to make three images of quasi-stellar object 1304+295 to find out if the object has sufficient ultraviolet flux levels for subsequent observations using the GHRS and FOS. All three observations were received in the OSS. No problems were noted. 1.3 Completed Two Sets of FOS/BL 3683 (Accretion Disk Mapping in Eclipsing Cataclysmic variables) The Faint Object Spectrograph (Rapid Mode) was used to make spectral observations of the cataclysmic variable star IP Pegasi during its eclipse with its accretion disk. The initial peakup failed due to a loss of lock that occurred just shortly before the dwell scan began. (See Note 2.1 and 2.4) FGS control was regained in time for the second peakup to execute, but the OSS could not say if the target was well placed in the aperture. The first FOS rapid mode observation was scheduled too late to catch the eclipse. Three subsequent spectral observations were reported to look okay. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.5 Completed GHRS/2 4601 (Jupiter's Hydrogen Lyman Alpha Auroral/Airglow Line Profile: Cycle 3 Medium) The High Resolution spectrograph was used to make spectral observations of the Hydrogen I 1216 A0 line in Jupiter's atmosphere at latitudes of +40, +50, +55, +60, and +70 degrees on the central meridian. This is looking of emissions from the auroral torus. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 4 Successful : 4 Losses of Lock : 1 Loss of Fine Lock, 148/0731Z, during Proposal FOS 3683, Reacquisition A-Okay (See Note 2.4 and 1.3). 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Note: Special target of opportunity observations will continue using the Faint Object Spectrograph to make spectral observations of cataclysmic contact variable star IP Pegasi during predicted eclipse periods. Proposal FOS #3683 (See Note 1.3 and 2.1) Observation times: 147/0715Z - 0945Z (Finished) 147/1830Z - 2100Z (Finished) 148/0545z - 0815Z (Finished) 149/2331z - 150/0200Z Wednesday May 26 ---------------- The Breathing Test was executed on the spacecraft on May 25-26. Preliminary indicators are that the test proceeded as planned. There were sessions of two data takes each. Each session included guide star acquisitions using fine lock, followed by parallel data takes involving FGS #2 as the astrometry FGS, and images from the FOC and WF/PC. Data analysis is now under way. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FGS 4725 (Scale Cycle 2 Calibration) 3.2 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 3.3 Complete Two Sets of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) 3.4 Complete Four Sets of PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) 3.5 Complete GHRS/2 4601 (Jupiter's Hydrogen Lyman Alpha Auroral/ Airglow Line Profiles: Cycle 3 Medium) 3.6 Complete FOC/96 4800 (The Stellar Content of Wolf - Rayet Galaxies (FOC Images)) 3.7 Complete WFC 4381 (HST Medium - Deep Survey Cycle 3) 3.8 Complete FGS 2859 (Linking HIPPARCOS to the Extragalactic Reference Frame Part 2 of 6) 3.9 Complete FGS 4031 (High Speed Astrometry - A search for Planetary Companions to Low Mass) 3.10 Complete FOC/96 4152 (FOC Observations to Monitor Distortion) 3.11 Complete FOC/96 4153 (FOC UV Throughput Monitoring) 3.12 Complete WFC/PC 4109 (WF/PC Cycle 2 Calibration: |Rapid Internal Monitor) 3.13 Complete PC 4929 (WF/PC cycle 2: Improved Dark Calibration) 3.14 Completed Two Sets of FOS/BL 3683 (Accretion Disk Mapping in Eclipsing Cataclysmic variables) 3.15 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.16 Complete GHRS/96 4569 (A Search for Microflaring Activity in Dwarf M Stars - Cycle 3 High) 3.17 Complete GHRS/2 3444 (Composition of Gas in Individual Interstellar Clouds: Cycle 2 Augmentation) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The DF-224 flight software installation is scheduled for June 9, 1993. The DF-224 FSW FRR meeting is scheduled for Friday June 4. /DMH -END-